OpenStudy (anonymous):

Write the equation of a line that passes through the point (1,-3) and is parallel to the line 7x + y = -5.

OpenStudy (anonymous):

If its parallel to the line y= -7x-5 it means that the slope is the same of that line. All you need to do now is to find the y- axis intersection. Now putting it into numbers: If two lines are parallel, the slope has the same value (-7) You also know that it needs to pass through point (1,-3) then: y = -7x +b -3 = -7*1 + b (point (1,-3) b= 4 then you equation is y = -7x + 4